# Fidel-TS: A High-Fidelity Benchmark for Multimodal Time Series Forecasting

### v2

@@ Abstract
The evaluation of time series forecasting models is hindered by a critical lack of high-quality benchmarks, leading to a potential illusion of progress. Existing datasets suffer from issues ranging from pre-training data contamination in the age of LLMs to the causal and description leakage prevalent in early multimodal designs. To address this, we formalize the core principles of high-fidelity benchmarking, focusing on data sourcing integrity, strict causal soundness, and structural clarity. We introduce Fidel-TS, a new large-scale benchmark built from the ground up on these principles by sourcing data from live APIs. Our extensive experiments validate this approach by exposing the critical biases and design limitations of prior benchmarks. Furthermore, we conclusively demonstrate that the causal relevance of textual information is the key factor in unlocking genuine performance gains in multimodal forecasting.
@@ 1 Introduction
The field of time series forecasting has seen relentless progress, with a proliferation of methods from classic statistical models (Box et al., 2015) and machine learning models (Chen & Guestrin, 2016) to sophisticated deep learning architectures (Hochreiter & Schmidhuber, 1997; Vaswani et al., 2017) and foundation models (Garza et al., 2023; Rasul et al., 2023). As models demonstrate increasingly strong capabilities, the field of time-series forecasting calls for evaluations that are broader, fairer, and more aligned with real-world scenarios. However, we argue that the benchmarks currently in use have not kept pace with the rapid advancements in model development.
@@
This weak foundation is rooted in legacy issues and compounded by new, misguided approaches. Classic unimodal benchmarks (Zhou et al., 2021; Wu et al., 2021; 2023) are often small in scale, outdated, and feature a vague composition of variables, making them scientifically unsound for evaluating modern models. More critically, the widespread availability of these datasets has created a significant, often unacknowledged, risk of pre-training contamination for large language models (LLMs) (Magar & Schwartz, 2022; Jiang et al., 2024), which may lead to the hallucination that LLMs excel in time series forecasting. As the community has pivoted to multimodal forecasting (Williams et al., 2024; Wang et al., 2024a; Xu et al., 2024a), first-generation benchmarks (Liu et al., 2024a) have perpetuated this contamination risk while introducing equally severe flaws. The common practice of retrospectively retrieving textual data introduces severe data leakage, including causal leakage (accessing information published after the event to be predicted) and description leakage (retrieving text that explicitly states the ground-truth value). Furthermore, relying on low-frequency data (e.g., monthly) makes these datasets too small to effectively train domain-specific numerical models and prevents them from providing sufficiently complex testing scenarios.
@@
To rectify these shortcomings and establish a more credible evaluation standard, we reconsider the essential characteristics of data encountered in real-world forecasting scenarios. Based on this review, we formalize core principles of high-fidelity benchmarking: (1) Data Sourcing Integrity, which mandates the use of real-time, authentication-protected API data streams to not only ensure recency and data sufficiency but also mitigate pre-training contamination; (2) Strict Causal Soundness, which incorporates only verifiably exogenous textual information, such as weather forecasts and scheduled maintenance, to prevent causal and description leakage; (3) Structural Clarity, which implements a clear demarcation between forecasting "Subjects" and data "Channels" to enable rigorously evaluating model generalization. Guided by these paradigms, we introduce Fidel-TS, a new benchmark designed from the ground up to embody these principles. Each dataset contains millions of fresh, high-frequency and leak-free data points from real-time APIs, features textual information covering every timestamp, and is built with a highly extensible architecture.
To demonstrate the value of our high-fidelity benchmark, we developed a comprehensive framework compatible with models of different modalities, architectures, and input. Our experiments reveal substantial differences in the performance of many state-of-the-art models on Fidel-TS compared to prior benchmarks, and further uncover the poor performance of LLMs on forecasting tasks on Fidel-TS, indicating the evaluation bias and design flaws of prior benchmarks. Leveraging the benchmark’s causally sound design, we investigate how different textual inputs affect multimodal forecasting. Our results show that performance gains hinge on the causal relevance of the text, underscoring causal alignment as an important factor for designing robust datasets and effective models. In summary, our contributions are threefold:
@@ 2 Related Works
2.1 Classic Unimodal Benchmarks
The evaluation of time series forecasting models has historically relied on a narrow set of classic unimodal benchmarks, typically including ETT(ETTh1, ETTh2, ETTm1, ETTm2), ECL (Zhou et al., 2021); Electricity, Weather, Traffic, ILI (Wu et al., 2021); M4 (Wu et al., 2023), etc. The scientific utility of classic benchmarks is compromised by three core issues rooted in their age and static nature:
@@ 2.1 Classic Unimodal Benchmarks
2.2 Novel Multimodal Benchmarks
@@
Second, the reliance on an Agent pipeline to retrieve web-based text introduces multiple, critical forms of data leakage. This method risks pre-training contamination, as static web content is likely already part of the training corpora for LLMs. It also leads to causal leakage, where retrieved documents like future financial reports contain information unavailable at prediction time, and description leakage, where text explicitly describes future trends, effectively revealing the ground truth. Furthermore, this approach fails to ensure strict temporal alignment between the textual data and its corresponding time series timestamp, increasing the risk of causal leakage issue.
@@
Third, it perpetuates the ambiguous variable structure of classic unimodal benchmarks. It also fails to distinguish between datasets comprising multiple, distinct variables from a single system versus those with a single variable type from multiple systems. Besides, in its multivariate datasets, it arbitrarily designates one channel as the "target," a design choice that oversimplifies the problem and ignores the equal importance, or potential interdependencies among variables.
Taken together, these pervasive issues of data scarcity, data leakage, and ambiguous structure across existing benchmarks result in a significant loss of real-world fidelity. This foundational weakness affects the reliable evaluation of all models and hinders credible progress. To rectify this, we argue for a new evaluation paradigm founded on a clear set of principles, which we introduce next.
3 Fidel-TS: A Benchmark Built on High-Fidelity Principles
3.1 Core Principles of High-Fidelity Design
The foundation of our benchmark is the exclusive use of continuously updated data streams from authentication-protected APIs, eschewing the static, web-scraped files common in other benchmarks. This single architectural choice directly addresses several requirements for realism.
First, it provides a stream of large quantity, fresh, high-frequency data, ensuring the benchmark remains current and can be used to evaluate models on fine-grained, sub-hourly tasks. This API-driven approach also guarantees precise temporal alignment across all modalities, a critical factor for accurate forecasting.
Second, this methodology provides a multi-layered defense against information leakage. We mitigate the risk of pre-training contamination through two primary mechanisms. First, by sourcing data directly from APIs rather than from static, widely-indexed public files, we reduce the likelihood that our benchmark is part of common training corpora. More critically, the recency of our data streams, which are continuously updated and extend well beyond the typical knowledge cutoffs, serves as a robust safeguard against direct dataset memorization. Furthermore, the strict chronological data delivery inherent to the API-driven approach naturally prevents the forms of causal leakage that arise from misaligned or retrospectively collected data.
With a foundation of fresh, well-aligned data, our next principle is to populate it with causally sound, context-rich information. To prevent causal leakage and description leakage, the benchmark exclusively incorporates verifiably independent, exogenous factors. For example, providing a future weather forecast (’Temperature will drop tomorrow’) is a valid causal input, whereas providing a description of the outcome (’Expect high power usage tomorrow due to cold wave’) would constitute description leakage. Our primary source of such information is weather data—a near-universal factor that influences most real-world systems while remaining causally independent of them. For relevant datasets, we also include planned operational signals (e.g., scheduled device maintenance), which represent another class of knowable future events. This rich context of time-series data and aligned textual descriptions provides a comprehensive basis for evaluating a model’s ability to reason about complex, real-world dynamics.
In this foundational version of Fidel-TS, we deliberately focus on weather as the primary textual modalities on most of our datasets. While other sources like news or economic reports offer richer linguistic diversity, they are notoriously difficult to align temporally and are fraught with the risk of the very causal and description leakage our principles are designed to eliminate. Our focus on these cleaner, verifiably exogenous sources is a direct reflection of our commitment to first establishing a gold-standard, leak-free baseline for causal multimodal evaluation.
Finally, the entire framework is designed for extensibility. By providing detailed metadata and documentation for all API sources, we enable researchers to easily integrate additional data streams—such as new economic indicators or social media trends. This mirrors the real-world requirement for forecasting systems to adapt and scale as new, relevant information becomes available, ensuring the long-term relevance and utility of the benchmark.
3.2 Data Curation Pipeline: From Raw Streams to a High-Fidelity Benchmark
Translating raw, real-time data streams into a structured benchmark requires a pipeline that preserves the nuances of real-world data. The following steps detail our process, with a full summary of the complete pipeline in Figure 1 and the resulting datasets in Table 1. Detailed descriptions of the datasets are provided in Appendix A.1.
@@ 3.2 Data Curation Pipeline: From Raw Streams to a High-Fidelity Benchmark
Figure 1: The Construction Pipeline of Fidel-TS. The process integrates raw time series from diverse subjects and channels with dynamic textual information, all from real-time API. The Imperfection Handling step interpolates short data gaps while converting long downtime into time-aligned ’Sensor Downtime’ events. This enriched data, combined with static metadata, forms a unified dataset that is systematically organized into Realistic Benchmark Settings: downsampled (-h), sampled via importance (-mini), observed (-obs, for in-domain evaluation), and hidden (-hid, for generalization evaluation).
@@
Real-world forecasting requires both foundational knowledge and real-time updates. To provide this, we integrate two distinct types of textual information. First, we establish the static context by sourcing background knowledge from the official API metadata: a general overview for the entire dataset, specific descriptions for each Subject, and detailed explanations for each Channel. Second, we incorporate dynamic, time-aligned information that reflects the evolving state of the system and its environment. Weather serves as a common and universally accessible source of this dynamic data, as it influences most real-world systems without being directly influenced by them, and up-to-date information is readily available from open APIs. For relevant datasets, we also support the use of control information extracted from system logs or scheduled events. All textual data, both static and dynamic, is precisely aligned with the time series data inherently. See Appendix A.2 for examples of our preprocessed time series and textual data.
@@
To test model performance under diverse real-world conditions, we create several dataset variations. First, to ensure a fair comparison across datasets with different native sampling frequencies, we provide an option to downsample high-frequency series to a uniform hourly level (denoted by the -h suffix). Second, to address the unique computational challenges of evaluating LLMs, we curate smaller subsets (denoted by the -mini suffix) constructed via importance sampling. By selecting samples from the full test set where unimodal models outperform multimodal ones and vice versa, we ensure both computational efficiency and fairness.
@@
A crucial aspect of real-world utility is a model’s ability to generalize to new scenarios it has not been explicitly trained on. Our Subject and Channel definitions naturally support this evaluation. For datasets with a large number of Subjects, we create distinct partitions to simulate a zero-shot learning environment. Subsets intended for standard training and evaluation are marked with the -obs (observed) suffix, while those held back for testing generalization on unseen subjects are marked with the -hid (hidden) suffix, providing a robust test of a model’s transfer learning capabilities. More details are provided in Appendix A.3.
@@ 4.1 Evaluation Framework
The high-fidelity nature of our Fidel-TS benchmark establishes a robust foundation for evaluation. By supplying rich, leakage-free data, it enables not only a critical re-assessment of traditional unimodal models but also a sound basis for evaluating modern multimodal models and LLMs. However, the forecasting model landscape has become increasingly heterogeneous, with vast differences in model scale, architectural backbones, and input-output definitions. This diversity highlights a pressing need for a universal framework that enables fair and convenient comparisons. Existing toolkits (such as LSTF-Linear (Zeng et al., 2023), TSLib (Wang et al., 2024b), MM-TSFLib (Liu et al., 2024a), TSFM-Bench (Li et al., 2025)) often tailored to specific model categories and lack the flexibility to accommodate this wide spectrum of models under a single, unified configuration.
To bridge the gap, we developed our Universal Cross-modal Evaluation Framework designed to connect our high-fidelity data with the diverse model ecosystem. It supports both raw text and pre-computed embedding vectors while rigorously enforcing the temporal alignment between textual data and time series. To cater to the varied nature of modern models, our framework not only provides standard PyTorch (Paszke et al., 2019) interface, but also integrates HuggingFace Transformers (Wolf et al., 2020) for foundation models, and leverages PyTorch Lightning (Falcon et al., 2019) to accelerate the training of complex multimodal models. Recognizing the unique requirements of LLMs, it also supports both local deployment via vLLM (Kwon et al., 2023) and remote API calls through a simple socket. This thoughtful engineering ensures that our framework serves as the essential bridge between the rich, multimodal data of Fidel-TS and the diverse array of models, enabling the fair and comprehensive experiments that follow. The detailed comparison between our framework and prior frameworks is in Appendix B.

### v3

# Fidel-TS: A High-Fidelity Multimodal Benchmark for Time Series Forecasting

@@ Abstract
The evaluation of time series forecasting models is hindered by a critical lack of high-quality benchmarks, leading to a potential illusion of progress. Existing datasets suffer from issues ranging from pre-training data contamination in the age of LLMs to the temporal and description leakage prevalent in early multimodal designs. To address this, we formalize the core principles of high-fidelity benchmarking, focusing on data sourcing integrity, leak-free and causally sound design, and structural clarity. We introduce Fidel-TS, a new large-scale benchmark built from the ground up on these principles by sourcing data from live APIs. Our experiments reveal the flaws of the previous benchmarks and the biases in model evaluation, providing new insights into multiple existing forecasting models and LLMs across various evaluation tasks. The code is available at https://github.com/VEWOXIC/Universal-Cross-Modal-Time-Series-Forecasting-Pipeline and the data is available at https://huggingface.co/collections/wxcai/fidel-ts-68ef3c7813c4187e2c88dedd
Machine Learning, ICML
@@ 1 Introduction
The field of time series forecasting has seen relentless progress, with a proliferation of methods from classic statistical models (Box et al., 2015) and machine learning models (Chen and Guestrin, 2016) to sophisticated deep learning architectures (Hochreiter and Schmidhuber, 1997; Vaswani et al., 2017) and foundation models (Garza et al., 2023; Rasul et al., 2023). As models demonstrate increasingly strong capabilities, the field of time-series forecasting calls for evaluations that are broader, fairer, and more aligned with real-world scenarios. However, we argue that the benchmarks currently in use have not kept pace with the rapid advancements in model development.
@@
This weak foundation is rooted in legacy issues and compounded by new, misguided approaches. Classic unimodal benchmarks (Zhou et al., 2021; Wu et al., 2021, 2023) are often small in scale, outdated, and feature a vague composition of variables, making them scientifically unsound for evaluating modern models. More critically, the widespread availability of these datasets has created a significant, often unacknowledged, risk of pre-training contamination for large language models (LLMs) (Magar and Schwartz, 2022; Jiang et al., 2024), which may lead to the hallucination that LLMs excel in time series forecasting. As the community has pivoted to multimodal forecasting (Williams et al., 2024; Wang et al., 2024a; Xu et al., 2024a), first-generation benchmarks (Liu et al., 2024a) have perpetuated this contamination risk while introducing equally severe flaws. The common practice of retrospectively retrieving textual data introduces severe data leakage, including temporal leakage (accessing information published after the event to be predicted) and description leakage (retrieving text that explicitly states the ground-truth value). Furthermore, relying on low-frequency data (e.g., monthly) makes these datasets too small to effectively train domain-specific numerical models and prevents them from providing sufficiently complex testing scenarios.
@@
To rectify these shortcomings and establish a more credible evaluation standard, we reconsider the essential characteristics of data encountered in real-world forecasting scenarios. Based on this review, we formalize core principles of high-fidelity benchmarking: (1) Data Sourcing Integrity, which mandates the use of real-time, authentication-protected API data streams to not only ensure recency and data sufficiency but also mitigate pre-training contamination; (2) Leak-free and Causally Sound Design, which incorporates only verifiably exogenous textual information, such as weather forecasts and scheduled maintenance, to prevent temporal and description leakage; (3) Structural Clarity, which implements a clear demarcation between forecasting "Subjects" and data "Channels" to enable rigorously evaluating model generalization. Guided by these paradigms, we introduce Fidel-TS, a new benchmark designed from the ground up to embody these principles. Each dataset contains millions of fresh, high-frequency and leak-free data points from real-time APIs, features textual information covering every timestamp, and is built with a highly extensible architecture.
To demonstrate the value of our high-fidelity benchmark, we developed a comprehensive framework compatible with models of diverse modalities and architectures. This assessment reveals performance gap compared to earlier reports, exposing the evaluation biases inherent in prior benchmarks. From this evaluation, we obtain three key findings. First, in unimodal forecasting, no model consistently dominates, and Time Series Foundation Models fail to exhibit the claimed robust zero-shot generalization. Second, multimodal gains are not inherent but depend on model architecture and the causal relevance of textual inputs. Third, although LLM performance correlates with general reasoning ability, even domain-finetuned variants remain unstable; under realistic, context-rich, and leakage-free long-term settings, LLMs suffer from severe precision and reliability issues and consistently underperform dedicated forecasting models.
In summary, our contributions are:
@@ 2 Related Works
2.1 Unimodal Benchmarks for Time Series Forecasting
The evaluation of time series forecasting models has historically relied on a narrow set of classic unimodal benchmarks, typically including ETT(ETTh1, ETTh2, ETTm1, ETTm2), ECL (Zhou et al., 2021); Electricity, Weather, Traffic, ILI (Wu et al., 2021); M4 (Wu et al., 2023), etc. The scientific utility of these benchmarks is compromised by three core issues rooted in their age and static nature:
@@ 2.1 Unimodal Benchmarks for Time Series Forecasting
2.2 Multimodal Benchmarks for Time Series Forecasting
@@
Second, the reliance on an Agent pipeline to retrieve web-based text introduces multiple, critical forms of data leakage. This method risks pre-training contamination, as static web content is likely already part of the training corpora for LLMs. It also leads to temporal leakage, where retrieved documents like future financial reports contain information unavailable at prediction time, and description leakage, where text explicitly describes future trends, effectively revealing the ground truth. Furthermore, this approach fails to ensure strict temporal alignment between the textual data and its corresponding time series timestamp, increasing the risk of temporal leakage issue.
@@
Third, it perpetuates the ambiguous variable structure of unimodal benchmarks. It also fails to distinguish between datasets comprising multiple, distinct variables from a single system versus those with a single variable type from multiple systems. Besides, in its multivariate datasets, it arbitrarily designates one channel as the "target," a design choice that oversimplifies the problem and ignores the equal importance, or potential interdependencies among variables.
@@
Taken together, these pervasive issues of data scarcity, data leakage, and ambiguous structure across existing benchmarks result in a significant loss of real-world fidelity, further impact on the evaluation of the models.
3 Fidel-TS: A Benchmark Built on High-Fidelity Principles
Figure 1: The Construction Pipeline of Fidel-TS. The workflow integrates raw data from the Target System Time Series API and the Dynamic Heterogeneous API. The Imperfection Handling step interpolates short data gaps while converting long downtime into time-aligned ’Sensor Downtime’. These downtime events and heterogeneous textual data are stored in the Dynamic Event Database together, strictly following the start and end times. All these events and time series, combined with static dataset metadata, forms a unified dataset that is systematically organized into Realistic Benchmark Settings: downsampled (-h), sampled via importance (-mini), observed (-obs, for in-domain evaluation), and hidden (-hid, for generalization evaluation).
3.1 Core Principles of High-Fidelity Design
The foundation of our benchmark is the exclusive use of continuously updated data streams from authentication-protected APIs, eschewing the static, web-scraped files common in other benchmarks. This single architectural choice directly addresses several requirements for realism. First, it provides a stream of large quantity, fresh, high-frequency data, ensuring the benchmark remains current and can be used to evaluate models on fine-grained, sub-hourly tasks. This API-driven approach also guarantees precise temporal alignment across all modalities, a critical factor for accurate forecasting. Second, this methodology provides a multi-layered defense against information leakage. We mitigate the risk of pre-training contamination through two primary mechanisms. First, by sourcing data directly from APIs rather than from static, widely-indexed public files, we reduce the likelihood that our benchmark is part of common training corpora. More critically, the recency of our data streams, which are continuously updated and extend well beyond the typical knowledge cutoffs, serves as a robust safeguard against direct dataset memorization.
With a foundation of fresh, well-aligned data, our next principle is to populate it with leak-free, causally sound and context-rich textual information. We deliberately focus on weather reports (sourced from weather APIs) as the primary textual modality for the majority of our datasets. First, weather is a variable that is widely recognized in the fields of photovoltaics, electricity, and transportation as a variable that can guide the prediction rather than merely serving as a historical covariate (Antonanzas et al., 2016) (Hong and Fan, 2016) (Polson and Sokolov, 2017). Unlike economic reports or news updates, weather forecasts can be known in advance. This verifiable availability prevents temporal leakage. Second, weather data effectively circumvents description leakage. Whereas some textual inputs might explicitly describe the target variable (e.g., "power usage will rise next week"), weather reports strictly describe environmental conditions without directly stating the outcome of the variable itself. Finally, weather is causally sound and verifiable: based on real-world common sense, weather is a universal driver that significantly influences system dynamics, yet it remains an exogenous factor independent of and unaffected by the systems it influences.
Finally, the entire framework is designed for extensibility. On one hand, researchers can obtain the latest data of the dataset from the real-time API at any time through our data construction process. On the other hand, researchers can easily use the flow to add additional time series and heterogeneous data and align them in time.
3.2 Data Curation Pipeline
Translating raw, real-time data streams into a structured benchmark requires a pipeline that preserves the nuances of real-world data. The following steps detail our process, with a full summary of the complete pipeline in Figure 1 and the resulting datasets in Table 1. More detailed background information of these datasets are provided in Appendix A.1.
@@ 3.2 Data Curation Pipeline
Real-world forecasting requires both foundational knowledge and real-time updates. To provide this, we integrate two distinct types of textual information. First, we establish the static context by sourcing background knowledge from the official API metadata: a general overview for the entire dataset, specific descriptions for each Subject, and detailed explanations for each Channel. Second, we incorporate dynamic, time-aligned information that reflects the evolving state of the system and its environment. This dynamic information, along with detected sensor downtime events in Imperfection Handling step, is consolidated into the Dynamic Event Database, which features storage indexed by start and end times, ensuring that the textual context is strictly aligned with the timestamps of the numerical time series. For all datasets, we provide weather reports for their respective locations. For relevant datasets, we also support the use of control information extracted from system logs or scheduled events. See Appendix A.2 for examples and visual analysis of our preprocessed time series and textual data.
@@
To test model performance under diverse real-world conditions, we create several dataset variations. First, to ensure a fair comparison across datasets with different native sampling frequencies, we provide an option to downsample high-frequency series to a uniform hourly level (denoted by the -h suffix). Second, to address the unique computational challenges of evaluating LLMs, we curate smaller subsets (denoted by the -mini suffix) constructed via importance sampling. By selecting samples from the full test set where unimodal models outperform multimodal ones and vice versa, we ensure both computational efficiency and fairness. More details of the algorithm are provided in Appendix A.3.
@@
A crucial aspect of real-world utility is a model’s ability to generalize to new scenarios it has not been explicitly trained on. Our Subject and Channel definitions naturally support this evaluation. For datasets with a large number of Subjects, we create distinct partitions to simulate a zero-shot learning environment. Subsets intended for standard training and evaluation are marked with the -obs (observed) suffix, while those held back for testing generalization on unseen subjects are marked with the -hid (hidden) suffix, providing a robust test of a model’s transfer learning capabilities. More details of the division are provided in Appendix A.4.
@@ 4.1 Evaluation Framework
The high-fidelity nature of our Fidel-TS benchmark establishes a robust foundation for evaluation. By supplying rich, leakage-free data, it enables not only a critical re-assessment of traditional unimodal models but also a sound basis for evaluating modern multimodal models and LLMs. However, the forecasting model landscape has become increasingly heterogeneous, with vast differences in model scale, architectural backbones, and input-output definitions. This diversity highlights a pressing need for a universal framework that enables fair and convenient comparisons.
Table 2: Comparison of time series forecasting frameworks across four key attributes
Attributes Benchmark Frameworks
LSTF- Linear TSLib TSFM- Bench MM- TSFLib Ours
Model Domain-Specific $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Foundation Models $\checkmark$ $\checkmark$
Structure Numerical $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
LLM-based $\checkmark$ $\checkmark$
LLM APIs $\checkmark$ $\checkmark$
Modality Unimodal $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Multimodal $\checkmark$ $\checkmark$
Setting Full-shot $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Zero-shot $\checkmark$ $\checkmark$
Existing toolkits (such as LSTF-Linear (Zeng et al., 2023), TSLib (Wang et al., 2024b), MM-TSFLib (Liu et al., 2024a), TSFM-Bench (Li et al., 2025)) often tailored to specific model categories and lack the flexibility to accommodate this wide spectrum of models under a single, unified configuration. To bridge the gap, we developed our Universal Cross-modal Evaluation Framework designed to connect our high-fidelity data with the diverse model ecosystem. The comparison in Table 2 illustrates the comprehensiveness of our benchmark across training paradigms, model types, modality support, and evaluation strategies. Our framework provides an integrated testing environment for diverse models and facilitates unified experimental configuration. More implementation details of our framework are provided in Appendix B.